OpenStudy (anonymous):

Find the sum of the first 62 terms of the arithmetic sequence 35,34, 33, 32........

OpenStudy (campbell_st):

ok the sum of an arithmetic sequence can be found using the formula \[S_{n} = \frac{n}{2}[2a + (n - 1)d]\] where n = number of terms a = 1st term d = common difference in your question you have n = 62 a = 35 d = -1 just substitute into the formula and evaluate

OpenStudy (anonymous):

I can't figure it out still.....

OpenStudy (anonymous):

like i did what you say but i got something that wasn't right

OpenStudy (campbell_st):

ok here it is \[S_{62} = \frac{62}{2}[ 2 \times 35 + (61 - 1) \times (-1)]\] doing it in bits you get \[S_{62} = 31[70 + 61 \times (-1)]\] this may help

OpenStudy (campbell_st):

what answer did you get... ?